1. You own a pole that is 80m long and a barn that is 40m long. You decide to pick up the pole and run with it through the barn. How fast must you run with the pole so that (for a moment, at least!) the pole will completely fit inside the barn? Give your answer as a fraction of the speed of light; for example, v=0.5c
Answer #1

Let the speed be v,
L = Lo * sqrt(1 - v^2/c^2)
Where
L = 40 m
Lo = 80 m

Substituting values,
40 = 80 * sqrt(1 - v^2/c^2)
(1/2)^2 = 1 - v^2/c^2
v^2/c^2 = 1 - 1/4
v = c*sqrt(3)/2
v = 0.866 c